OpenStudy (anonymous):

Use a sum or difference formula (not a calculator or computer!) to determine the value of cos(7π/12). a. (√2)/4+(√6)/4 b. 5/6 c. (√2)/4-(√6)/4 d. (√3)/4

OpenStudy (anonymous):

\[\frac{\pi}{3}+\frac{\pi}{4}=\frac{7\pi}{12}\]

OpenStudy (anonymous):

\[\cos(\alpha +\beta)=\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)\] put \(\alpha =\frac{\pi}{3}, \beta=\frac{\pi}{4}\)

OpenStudy (anonymous):

Would it be C?

OpenStudy (anonymous):

yeah as usual it is C

OpenStudy (anonymous):

I used a calculator, would the answer be different if I didn't use one?